Suppose the temperature of a COVID positive patient is recorded as Tbody = 40°C.



Determine the wavelength at which the radiation emitted from the body reaches its peak.

1
Expert's answer
2022-02-04T07:40:44-0500

Solution

Temperature

T=40°C=313K

Using wein law

"\lambda=\frac{b}{T}\\=\frac{ 2.898 \times 10^{ -3 } }{313}\\=92.59A°"